OpenStudy (anonymous):

what polynomial identity should be used to prove that 35=8+27?

OpenStudy (mathmath333):

\(\large\tt \ \begin{align} \color{black}{8+27\\~\\ =2^3+3^3\\~\\ =(2+3)(2^2-2\times 3+3^2)\\~\\ =5\times 7\\~\\ =35}\end{align}\)